# In-Context Learning Dynamics with Random Binary Sequences

### v1

@@ 3 Algorithmic and Subjective Randomness
Next, we describe the domain of randomness that we explore in this work, and provide relevant background from computational cognitive science. To build intuition, consider: which sequence of coin flips is more random, $\mathtt{TTHTHHTH}$ or $\mathtt{HTTTTTTT}$? While the odds of a fair coin producing either sequence are the same (1 in 256), most people say that the first sequence “looks more random”. In a bias termed the Gambler’s Fallacy, people reliably perceive binary sequences with long streaks of one value as less random, and judge binary sequences with higher-than-chance alternation rates as being “more random” than truly random sequences (Falk & Konold, 1997). In addition to not using simple objective measures when judging the randomness of sequences, people also have a hard time generating sequences that are truly random  (Oskarsson et al., 2009; Griffiths et al., 2018; Gronchi & Sloman, 2021; Meyniel et al., 2016; Planton et al., 2021).
@@
Cognitive scientists studying Subjective Randomness model how people perceive randomness, or generate data that is subjectively random but algorithmically pseudo-random. One way to study this is to ask people whether a given data sequence was more likely to be generated by a Random process or a Non-Random process (Fig. 3). While the posterior distribution of all non-random processes is infinite, highly irregular, and includes every possible computable function, the problem of estimating this distribution can be simplified to finding the single most probable algorithm to approximate the full hypothesis space. If the hypotheses are programs which generate the data, a natural prior $p(h)$ is to assign higher probabilities to programs with shorter description lengths, or lower complexity. For example (Fig. 3), $\mathtt{HTTTTTTT}$ might be described with a shorter program—one H followed by sever Ts – than $\mathtt{HTTHTHHT}$, which cannot be compressed—H followed by two Ts, then one H, then one T, then two Hs, then one T. This optimization problem, searching for the shortest program consistent with data, is equivalent to computing the Kolmogorov complexity of a sequence $K(x)$ (Li et al., 1997) and has motivated the use of “simplicity priors” in a number of domains in computational cognitive science Chater & Vitányi (2003); Goodman et al. (2008); Piantadosi et al. (2012).
@@
Following previous work (Griffiths & Tenenbaum, 2003; Griffiths et al., 2018), here we define subjective randomness of a sequence as the ratio of likelihood of that sequence under a random versus non-random model, i.e., $\mathtt{randomness}(x)=\log P(x|\mathtt{random})\ -\ \log P(x|\mathtt{non}\text{-}\mathtt{random})$. The likelihood given a truly random Bernoulli process will be $p(x|\mathtt{random})=2^{-|x|}$, since each of $2^{|x|}$ binary sequences of length $|x|$ has equal probability. The non-random likelihood $p(x|\mathtt{non}\text{-}\mathtt{random})=2^{-K(x)}$ denotes the probability of the minimal description length program that generates $x$, equivalent to Bayesian model selection: $P(x|\mathtt{non}\text{-}\mathtt{random})=\max_{h\in\mathcal{H}}p(x|h)\ p(h)$. In theory, the space $\mathcal{H}$ of non-random models encompasses every possible computation; in this work, we study a small subset of this space that includes formal languages and probabilistic models inspired by psychological models of human concept learning and subjective randomness (Griffiths et al., 2018; Nickerson, 2002; Hahn & Warren, 2009). For further background, see Appendix H.
@@ 4 Experiments
Figure 4: Prompt templates and GPT-3.5 Randomness Judgments   (Left) Prompt templates used for the Randomness Generation (Top) and Judgment (Bottom) tasks. $\{\mathtt{sequence}\}$ is substituted with a list of context flips $x$, e.g. $\mathtt{Heads,\,Tails,\,Tails,\,Heads}$. $\{\mathtt{p}\}$ is substituted with the probability of Tails, and $\{\mathtt{1-p}\}$ with the probability of Heads. (Right) In Randomness Judgment tasks, we observe sharp transitions in the predictive distribution $p(y=\mathtt{random}|x)$ for GPT-3.5, from high-confidence in $x$ being generated by a random process, to high confidence in a non-random algorithm, as additional data $x$ is provided from the same concept.
Randomness Generation and Judgment Tasks
@@
We compare LLM-generated sequences to a ground truth “random” Bernoulli distribution with the same mean ($\mu=\overline{y}_{\text{LLM}}$), to a simple memory-constrained probabilistic model, and to Markov chains fit to model-generated data $y$. The Gambler’s Fallacy is the human bias to avoid long runs of a single value (‘Heads’ or ‘Tails’) consecutively in a row (Falk & Konold, 1997; Nickerson, 2002). Hahn & Warren (2009) theorize that the Gambler’s Fallacy emerges as a consequence of human memory limitations, where ‘seeming biases reflect the subjective experience of a finite data stream for an agent with a limited short-term memory capacity’. We formalize this as a simple Window Average model, which tends towards a specific probability $p$ as a function of the last $w$ flips: $p(y|x)=\max(0,\min(1,2p-\overline{x}_{t-w\ldots t}))$
Sub-Sequence Memorization and Complexity Metrics
@@
Bender et al. (2021) raise the question of whether LLMs are ‘stochastic parrots’ that simply copy data from the training set. We test the extent to which LLMs generate pseudo-random sequences by arranging memorized patterns of sub-sequences, rather than according to a more complex algorithm as our Window Average model represents. To measure memorization, we look at the distribution of unique sub-sequences in $y$. If an LLM is repeating common patterns across outputs, potentially memorized from the training data, this should be apparent in the distribution over length K sub-sequences. Since there are deep theoretical connections between complexity and randomness (Chaitin, 1977; 1990), we also consider the complexity of GPT-produced sequences. Compression is a metric of information content, and thus of redundancy over irreducible complexity (MacKay, 2003; Delétang et al., 2023), and neural language models have been shown to prefer generating low complexity sequences (Goldblum et al., 2023). As approximations of sequence complexity, we evaluate the distribution of Gzip-compressed file sizes (Jiang et al., 2022) and inter-sequence Levenshtein distances (Levenshtein et al., 1966).
Refer to captionRefer to captionRefer to caption
Figure 5: GPT-3.5 generates pseudo-random binary sequences that deviate from a Bernoulli process. (Left) Empirical conditional probabilities for a third-order Markov Chain fit to sequences $y$ generated by GPT-3.5 text-davinci-003, a Bernoulli process centered at the mean of GPT sequence $\overline{y}$, and our Window Average model ($w=5$). In the simulated Bernoulli process, edges are fairly uniform; the conditional probabilities for GPT-3.5 and the Window Average model demonstrate a similar non-uniform bias. (Right) Running averages for flip sequences generated by each model, where 00 denotes ‘$\mathtt{Heads}$’ and $1$ denotes ‘$\mathtt{Tails}$’. Compared to a Bernoulli process (top), sequences generating using GPT (middle) and those of our Window Average model (bottom) stay closer to the mean, repeating the same patterns more often.
Formal Language Learning Metrics
@@
In our Formal Language Learning analysis, $x$ is a subset of regular expression repetitions of short token sequences such as $x\in(\texttt{011})^{n}$, where longer sequences $x$ correspond to larger $n$. This enables us to systematically investigate in-context learning of formal languages, as $|x|$ corresponds to the amount of data for inducing the correct program (e.g. $(\texttt{011})^{n}$) from Bernoulli processes, as well as the space of all possible algorithms. A Markov chain of sufficiently high order $k$ can implicitly represent a regular language in its weights. However, the number of parameters required is exponential in Markov order $k$, and weights are difficult to interpret. For example, a 3-order Markov chain fit to data $x=(\texttt{011})^{n}$ will represent $(\texttt{011})^{n}$ corresponds to high weights in the conditional probability distribution for $p(\texttt{0}|\texttt{11})$, $p(\texttt{1}|\texttt{01})$, $p(\texttt{1}|\texttt{10})$, and low values for other conditional probabilities. In randomness Judgment tasks, we assess formal concept learning by the dynamics of $p(y=\mathtt{random}|x=C^{|x|})$ as a function of $|x|$. In randomness Generation tasks, we asses concept learning according to the language model predictive distribution $p(y|x)$ over output sequences, inferred from next-token generation data: $p(y_{0\ldots T}|x)=p(y_{0}|x)\prod_{t}^{T}p(y_{t}|y_{0,\ldots t-1},x)$. Given a space of possible outputs $y$ with length $d$, $y\in\{0,1\}^{d}$, we estimate $p(y|x)$ by enumerating all $y$ up to some depth $d$, and computing $\hat{p}(y_{d}|x,y_{1,\ldots,d-1})=\frac{1}{N}\sum_{i}^{N}(y_{d}==1)^{(i)}$ as the fraction of $N$ responses that are “Tails”. We use token-wise log probabilities $p(y_{t}|y_{0\ldots t-1})$ from the OpenAI API where available for cost efficiency and since this is equivalent to repeated token samples. The probability table for $p(y_{t}|y_{0,\ldots t-1})$ is a weighted binary tree with depth $d$, where edges represent the next-token probability $p(y_{t}|y_{t-1})$, and paths represent the probability of a sequence of tokens. We estimate the predictive probability $p(y_{t}\in C|x,y_{0\ldots,t-1})$ assigned to a given regular language by computing the total probability mass for all paths in $y_{0\ldots d}$ that exactly match $C$. For example, with $C=(\texttt{011})^{n}$, there will be 3 paths $y_{0\ldots d}$ that exactly match $C$, out of $2^{d}$ possible unique paths.
@@ 5.1 Subjectively Random Sequence Generation
Figure 6: GPT-3.5 shows a Gambler’s fallacy bias of avoiding long runs. (Top) Distribution of mean values of flip sequences ($\mu=\frac{1}{T}\sum_{t}y_{t}$) generated by GPT-3.5 (text-davinci-003) with the specified $p(Tails)$, compared with a Bernoulli process and our Window Average model with the same mean as the GPT-3.5 flips. Flips generated by GPT approximately follow the expected mean $p(Tails)$, but have lower variance than a Bernoulli distribution. (Bottom) Length of the longest run for each sequence, where a run is a sub-sequence of the same value repeating. In this case, we see a clear bias in GPT-3.5 to avoid long runs, with a similar pattern across all values of $p(Tails)$ despite the x-axis changing in scale.
In the GPT3.5 and higher (‘InstructGPT’) models — text-davinci-003, ChatGPT (gpt-3.5-turbo, gpt-3.5-turbo-instruct) and GPT-4 — we find an emergent behavior of generating seemingly random binary sequences (Fig. 5). This behavior is controllable, where different $p(\mathtt{Tails})$ values leads to different means of generated sequences $\overline{y}$. However, the distribution of sequence means, as well as the distribution of the length of the longest runs for each sequence, deviate significantly from a Bernoulli distribution centered at $\overline{y}$, analogous to the Gambler’s Fallacy bias in humans (see Fig. 6). Our Window Average model with a window size of $w=5$ partly explains both biases, matching GPT-generated sequences more closely than a Bernoulli distribution.
@@ 5.2 Sub-Sequence Memorization and Complexity
Figure 7: GPT repeats specific sub-sequences more often than chance. We find that GPT-3.5 repeats certain sub-sequences more often than is predicted by our Window Average model. While the Window Average model (green) generates fewer unique sub-sequences than a Bernoulli process (red), this does not account for the bias in GPT-3.5 (text-davinci-003, in blue) to repeat many of the same sub-sequences. This disparity increases with longer sub-sequences (see Fig. 20 in Appendix).
@@
We find significant differences between the distributions of sub-sequences for GPT-3.5 -generated sequences and sequences sampled from a Bernoulli distribution(see Fig. 7). This difference is partly accounted for with a Window Average model with a window size $w=5$, although GPT repeats certain longer sub-sequences, for example length-20 sub-sequences, that are far longer than 5. However, the majority of sub-sequences have very low frequency, and though further experiments would be required to conclude that all sub-sequences are not memorized from training data, it seems unlikely that these were in the training set, since we find thousands of unique length-k (with varying k) sub-sequences generated at various values of $P(\mathtt{Tails})$. This indicates that GPT-3.5 combines dynamic, subjectively random sequence generation with distribution-matched memorization. We also note that with a specification of $P(Tails)=50\%$, but not $49\%$, $51\%$ or other values, sequences $y$ generated by GPT-3.5+ are dominated by repeating ‘Heads, Tails, Heads, Tails, …’. This pattern is consistent across variations of the prompts listed in Fig. 4, including specifying ‘fair’ or ‘unweighted’ instead of a ‘weighted coin’, and produces a visible kink in many cross-$p(\mathtt{Tails})$ metrics (Fig. 232219). For this reason, in Fig. 6 we show results for $P(Tails)=51\%$.
@@
Across three metrics of sequence complexity — number unique sub-sequences, Gzip file size, and inter-sequence Levenshtein distance (see Fig. 2322 in Appendix) — we find that GPT-3.5+ models, with the exception of ChatGPT, generate low complexity sequences, showing that structure is repeated across sequences and supporting Goldblum et al. (2023); Delétang et al. (2023). By the metrics of mean Levenshtein distance and number of unique sub-sequences, ChatGPT generates higher complexity sequences than chance. We speculate that this phenomenon might explained by a cognitive model that avoids sampling with replacement.
Figure 8: With enough context, GPT-3.5 learns simple formal language concepts, transitioning from generating pseudo-random numbers to only generating values from the concept (Left) GPT-3.5 next-token predictive distribution $p(y|x)$ visualized as a binary tree, where red arrows correspond to Heads, blue arrows to Tails, and nodes matching the target concept $C=(011)^{n}$ are green. $p(y|x)$ changes with varying $|x|$, here $|x|=39$ and next-token predictions strongly follow the concept $C$. Also see Fig 1112 in Appendix. (Right) In-context learning dynamics for simple formal languages $x=(\texttt{HTH})^{n}$ (010) and $x=(\texttt{HTT})^{n}$ (011) as a function of context length $n=|x|$. Prediction accuracy computed as the total probability mass assigned to valid continuations of the formal language $x$, as a function of prediction depth $d=|y|$ and context length $|x|$. Curves shown are for $d=6$, where only 3 out of 64 total paths $y$ match concept $C$. Solid lines correspond to text-davinci-003 and dashed lines to gpt-3.5-turbo-instruct; note that learning curves for 010 and 011 flip between the two models. Also see Fig. 910 in Appendix.

### v3

@@ 3 Algorithmic and Subjective Randomness
Next, we describe the domain of randomness explored in this work, and provide relevant background from computational cognitive science. To build intuition, consider: which sequence of coin flips is more random, $\mathtt{TTHTHHTH}$ or $\mathtt{HTTTTTTT}$? While the odds of a fair coin producing either sequence are the same (1 in 256), most people say that the first sequence “looks more random”. In a bias termed the Gambler’s Fallacy, people reliably perceive binary sequences with long streaks of one value as ‘less random’, and judge binary sequences with higher-than-chance alternation rates as ‘more random’ than truly random sequences (Falk & Konold, 1997). In addition to not using simple objective measures when judging the randomness of sequences, people also have a hard time generating sequences that are truly random (Oskarsson et al., 2009; Griffiths et al., 2018; Gronchi & Sloman, 2021; Meyniel et al., 2016; Planton et al., 2021).
@@
Cognitive scientists studying Subjective Randomness model how people perceive randomness, or generate data that is subjectively random but algorithmically non-random, and have related this problem to algorithmic information theoretic definitions of randomness. One way to study this is to ask people whether a given data sequence was more likely to be generated by a Random process or a Non-Random process (Fig. 3). $\mathtt{HTTTTTTT}$ might be described with a shorter program—one H followed by seven Ts – than $\mathtt{HTTHTHHT}$, which cannot be compressed as well—H followed by two Ts, then one H, then one T, then two Hs, then one T. Viewing this task as program induction, a natural prior $p(h)$ over program hypotheses assigns higher weights to programs with shorter description lengths, or lower complexity. Similar to our description of model selection, or hypothesis search, compared to model averaging, the space of non-random programs can be simplified to a search for the single shortest program consistent with the data $x$. This search is equivalent to computing the Kolmogorov complexity of a sequence $K(x)$ (Li et al., 1997) and has motivated the use of “simplicity priors” in a number of domains in computational cognitive science Chater & Vitányi (2003); Goodman et al. (2008); Piantadosi et al. (2012).
@@
As in Griffiths & Tenenbaum (2003); Griffiths et al. (2018), we define subjective randomness of a sequence as the ratio of likelihood of that sequence under a random versus non-random model, i.e., $\mathtt{randomness}(x)=\log P(x|\mathtt{random})\ -\ \log P(x|\mathtt{non}\text{-}\mathtt{random})$. The likelihood given a truly random Bernoulli process is $p(x|\mathtt{random})=2^{-|x|}$, since each of $2^{|x|}$ binary sequences of length $|x|$ has equal probability. The non-random likelihood $p(x|\mathtt{non}\text{-}\mathtt{random})=2^{-K(x)}$ is the probability of the minimal description length program that generates $x$, equivalent to Bayesian model selection: $P(x|\mathtt{non}\text{-}\mathtt{random})=\max_{h\in\mathcal{H}}p(x|h)\ p(h)$. See App. K for more background. In theory, the space $\mathcal{H}$ of non-random models encompasses every possible computation; in this work, we study a small subset of this space that includes formal languages and probabilistic models inspired by psychological models of human concept learning and subjective randomness (Griffiths et al., 2018; Nickerson, 2002; Hahn & Warren, 2009). Specifically, we consider Bernoulli processes, regular languages, Markov chains, and a simple memory-constrained probabilistic model as candidates for $\mathcal{H}$.
@@ 4 Experiments
Our experiments aim to evaluate how LLMs operate in the domain of Subjective Randomness. Specifically, our goal is to understand whether ICL and text generation dynamics suggest model selection, or hypothesis search, over a discrete concept space. We first evaluate whether there is an emergent algorithm for generating subjectively random sequences, and if so, whether it is similar to cognitive models of human subjective randomness (Fig. 6,  7). We use open-ended text generation experiments (Fig. 4, Left) and analyze binary sequences produced by LLMs, and compare LLM sequences to a Bernoulli distribution and our cognitively inspired (Falk & Konold, 1997; Hahn & Warren, 2009) Window Average model. Second, we evaluate whether ICL dynamics in searching over random and non-random hypotheses demonstrate ICL operating as model averaging, where learning is smooth and continuous, or as model selection, or hypothesis search, where sudden, dramatic changes can occur during learning.
Figure 4: Prompt templates and GPT-3.5 Randomness Judgments   (Left) Prompt templates used for the Randomness Generation (top) and Judgment (bottom) tasks. $\{\mathtt{sequence}\}$ is substituted with a list of context flips $x$, e.g. $\mathtt{Heads,\,Tails,\,Tails,\,Heads}$. $\{\mathtt{p}\}$ is substituted with the probability of Tails, and $\{\mathtt{1-p}\}$ with the probability of Heads. (Right) In Randomness Judgment tasks, we observe sharp transitions in the predictive distribution $p(y=\mathtt{random}|x)$ for GPT-3.5, from high-confidence in $x$ being generated by a random process, to high confidence in a non-random algorithm, as additional data $x$ is provided from the same concept. See App. G for additional concepts.
@@
Subjective Randomness Models  We compare LLM-generated sequences to a ground truth “random” Bernoulli distribution with the same mean ($\mu=\overline{y}_{\text{LLM}}$), to a simple memory-constrained probabilistic model, and to Markov chains fit to model-generated data $y$. The Gambler’s Fallacy is the human bias to avoid long runs of a single value (‘Heads’ or ‘Tails’) consecutively in a row (Falk & Konold, 1997; Nickerson, 2002). Hahn & Warren (2009) theorize that the Gambler’s Fallacy emerges as a consequence of human memory limitations, where ‘seeming biases reflect the subjective experience of a finite data stream for an agent with a limited short-term memory capacity’. We formalize this as a simple Window Average model, which tends towards a specific probability $p$ as a function of the last $w$ flips: $p(y|x)=\max(0,\min(1,2p-\overline{x}_{t-w,\ldots,t}))$
Figure 5: With enough context, GPT-3.5 learns simple formal language concepts, transitioning from generating subjectively random sequences to only generating values matching the concept (Left) GPT-3.5 next-token predictive distribution $p(y|x)$ visualized as a binary tree, where red arrows correspond to Heads (00), blue arrows to Tails ($1$), and nodes matching the target concept $C=(011)^{n}$ are dark green. $p(y|x)$ changes with varying $|x|$, here $|x|=39$ and next-token predictions strongly follow the concept $C$. Also see Fig 1516 in Appendix. (Right) In-context learning dynamics for simple formal languages $x=(\texttt{HTH})^{n}$ (010) and $x=(\texttt{HTT})^{n}$ (011) as a function of context length $n=|x|$. Prediction accuracy computed as the total probability mass of $p(y|x)$ assigned to valid continuations of the formal language $x$, as a function of prediction depth $d=|y|$ and context length $|x|$. Curves shown are for depth $d=6$, where only 3 out of 64 possible length-6 binary sequences $y$ match concept $C$. Note: bottom model is gpt-3.5-turbo-instruct-0914. Also see App. F.
@@
Sub-Sequence Memorization and Complexity MetricsBender et al. (2021) raise the question of whether LLMs are ‘stochastic parrots’ that simply copy data from the training set. We test the extent to which LLMs generate subjectively random sequences by arranging memorized patterns of sub-sequences, rather than according to a more complex algorithm as our Window Average model represents. To measure memorization, we look at the distribution of unique sub-sequences in $y$. If an LLM is repeating common patterns across outputs, potentially memorized from the training data, this should be apparent in the distribution over length K sub-sequences. Since there are deep theoretical connections between complexity and randomness (Chaitin, 1977; 1990), we also consider the complexity of GPT-produced sequences. Compression is a metric of information content, and thus of redundancy over irreducible complexity (MacKay, 2003; Delétang et al., 2023), and neural language models have been shown to prefer generating low complexity sequences (Goldblum et al., 2023). As approximations of sequence complexity, we evaluate the distribution of Gzip-compressed file sizes (Jiang et al., 2022) and inter-sequence Levenshtein distances (Levenshtein et al., 1966).
@@
Formal Language Learning Metrics  In our Formal Language Learning analysis, $x$ is a subset of regular expression repetitions of short token sequences such as $x\in(\texttt{011})^{n}$, where longer sequences $x$ correspond to larger $n$. This enables us to systematically investigate in-context learning of formal languages, as $|x|$ corresponds to the amount of data for inducing the correct program (e.g. $(\texttt{011})^{n}$) from Bernoulli processes, as well as the space of all possible algorithms. A Markov chain of sufficiently high order $k$ can implicitly represent a regular language in its weights (see Fig. 6, Right). However, the number of parameters required is exponential in Markov order $k$, and weights are difficult to interpret. For example, a 3-order Markov chain fit to data $x=(\texttt{011})^{n}$ will represent $(\texttt{011})^{n}$ corresponds to high weights in the conditional probability distribution for $p(\texttt{0}|\texttt{11})$, $p(\texttt{1}|\texttt{01})$, $p(\texttt{1}|\texttt{10})$, and low values for other conditional probabilities. To match the goal of the Generation task, we use a variation of the prompt in Fig. 4: ‘Generate a sequence of 1000 samples that may be from a fair coin with no correlation, or from some non-random algorithm’
In Randomness Judgment tasks, we assess formal concept learning dynamics by changes in the predictive distribution $p(y=\mathtt{random}|x=C^{|x|})$, i.e. the probability of the final token being ‘Random’ vs. ‘Non’, as a function of context length $|x|$. The method for measuring concept learning dynamics in Randomness Generation tasks requires computing $p(y|x)$ over possible binary flip sequences $y$, instead of $y$ being a single token as with Judgment tasks. In binary sequence Generation tasks, we estimate the predictive probability $p(y_{t}\in C|x,y_{0\ldots,t-1})$ assigned to a given regular language concept $C$ by computing the total probability mass of $p(y|x)$ for all unique token sequences $y_{0\ldots d}$ up to some depth $d$, such that the $y$ exactly matches a permutation of the concept. If tokens in $x$ and $y$ have two possible values, 00 or $1$, the probability table for $p(y_{t}|y_{0,\ldots,t-1})$ is a weighted binary tree (see Fig. 5, Left) with depth $d$, where edges represent the next-token probability $p(y_{t}|y_{t-1})$, and paths represent the probability of a sequence of tokens. For example, with $C=(\texttt{011})^{n}$, there will be 3 paths $y_{0\ldots d}$ corresponding to the permutations $(\texttt{011})^{n},(\texttt{101})^{n},(\texttt{110})^{n}$, out of $2^{d}$ possible sequences $y$, and so: $p(y\in(\texttt{011})^{n}\ |\ x)=p(y=\text{`0, 1, 1, 0, 1, 1, \ldots'}\ |\ x)+p(y=\text{`1, 1, 0, 1, 1, 0, \ldots'}\ |\ x)+p(y=\text{`1, 0, 1, 1, 0, 1, \ldots'}\ |\ x)$ We use the notation $p(y\in C\ |\ x)$ in Fig. 5 (Right) to refer to this predictive probability of $y$ matching a concept $C$. In both Judgment and Generation tasks for formal concept learning, we use next-token prediction probabilities $p(y_{t}|y_{0,\ldots,t-1},x)$ directly from the OpenAI API, for cost efficiency and since this is equivalent to repeated token samples.
@@ 5.1 Subjectively Random Sequence Generation
Refer to caption    Refer to captionRefer to captionRefer to caption
Figure 6: GPT-3.5 generates subjectively random binary sequences that deviate from a Bernoulli process. (Left) Running averages for flip sequences generated by each model, where 00 denotes ‘$\mathtt{Heads}$’ and $1$ denotes ‘$\mathtt{Tails}$’. Compared to a Bernoulli process (top), sequences generating using GPT (middle) and those of our Window Average model (bottom) stay closer to the mean, repeating the same patterns more often. (Right) Empirical conditional probabilities for a third-order Markov Chain fit to sequences $y$ generated by GPT-3.5 text-davinci-003, a Bernoulli process centered at the mean of GPT sequence $\overline{y}$, and our Window Average model ($w=5$). In the simulated Bernoulli process, edges are fairly uniform; edges for GPT-3.5 and the Window Average model are non-uniform, but similar to one another.
In the GPT3.5 and higher models—InstructGPT (text-davinci-003), ChatGPT (gpt-3.5-turbo) and GPT-4—we find an emergent behavior of generating seemingly random binary sequences (Fig. 6). This behavior is controllable, where different $p(\mathtt{Tails})$ values lead to different means of generated sequences $\overline{y}$. However, the distribution of sequence means, as well as the distribution of the length of the longest runs for each sequence, deviate significantly from a Bernoulli distribution centered at $\overline{y}$, analogous to the Gambler’s Fallacy bias in humans (see Fig. 7). Our Window Average model with a window size of $w=5$ partly explains both biases, matching GPT-generated sequences more closely than a Bernoulli distribution. This result demonstrates that ICL enables selection of latent concepts in an LLM that operate as algorithms acting on data. Our next experiment, on sub-sequence memorization, aims to further characterize this algorithm and address the alternative hypothesis that this behavior is generated by memorizing pre-training data.
@@
Figure 7: GPT-3.5 shows a Gambler’s fallacy bias of avoiding long runs. (Top) Distribution of mean values of flip sequences ($\mu=\frac{1}{T}\sum_{t}y_{t}$) generated by GPT-3.5 (text-davinci-003) with the specified $p(\mathtt{Tails})$, compared with a Bernoulli process and our Window Average model with the same mean as the GPT-3.5 flips. Flips generated by GPT approximately follow the expected mean $p(\mathtt{Tails})$, but have lower variance than a Bernoulli distribution. (Bottom) Length of the longest run for each sequence, where a run is a sub-sequence of the same value repeating. In this case, we see a clear bias in GPT-3.5 to avoid long runs, with a similar pattern across all values of $p(\mathtt{Tails})$ despite the x-axis changing in scale.
@@ 5.2 Sub-Sequence Memorization and Complexity
We find significant differences between the distributions of sub-sequences for GPT-3.5 -generated sequences and sequences sampled from a Bernoulli distribution(see Fig. 8). This difference is partly accounted for with a Window Average model with a window size $w=5$, although GPT repeats certain longer sub-sequences, for example length-20 sub-sequences, that are far longer than 5. However, the majority of sub-sequences have very low frequency, and though further experiments would be required to conclude that all sub-sequences are not memorized from training data, it seems unlikely that these were in the training set, since we find thousands of unique length-k (with varying k) sub-sequences generated at various values of $P(\mathtt{Tails})$. This indicates that GPT-3.5 combines dynamic, subjectively random sequence generation with distribution-matched memorization. This suggests that ICL is indeed eliciting complex latent algorithms in subjective randomness tasks, and that this behavior cannot be easily attributed to memorizing sequences from the training distribution.
@@
Note that with a specification of $p(\mathtt{Tails})=50\%$, but not $49\%$, $51\%$ or other values, sequences $y$ generated by GPT-3.5+ are dominated by repeating $\mathtt{Heads},\mathtt{Tails},\mathtt{Heads},\mathtt{Tails},\dots$. This pattern is consistent across variations of the prompts listed in Fig. 4, including specifying ‘fair’ or ‘unweighted’ instead of a ‘weighted coin’, and produces a visible kink in many cross-$p(\mathtt{Tails})$ metrics (Fig. 252829). Due to this, in Fig. 7 we show results for $p(\mathtt{Tails})=51\%$. Across three metrics of sequence complexity—number unique sub-sequences, Gzip file size, and inter-sequence Levenshtein distance (see Fig. 2928 in Appendix)—we find that GPT-3.5+ models, with the exception of ChatGPT, generate low complexity sequences, showing that structure is repeated across sequences and supporting Goldblum et al. (2023); Delétang et al. (2023). For mean Levenshtein distance and number of unique sub-sequences, ChatGPT generates higher complexity sequences than chance. We speculate that this phenomenon might be explained in future work by a simple cognitive model: if the LLM makes a sequence more subjectively random by avoiding repeating sub-sequences, like a human might, then it will produce sequences with higher-than-chance complexity overall.